OpenStudy (beth12345):

how to solve for |2x|+3x=10 using algebra?

OpenStudy (jamesj):

If x > 0, then |2x| = 2x, hence the equation is 2x + 3x = 10 5x = 10 x = 2 As 2 > 0, this is consistent with the assumption x > 0 and therefore x = 2 is a solution. Now analyze the case x < 0

OpenStudy (beth12345):

would the x<0 part be like -(2x)<0, making the question -2x=-3x+10 and become x =-2?

OpenStudy (jamesj):

If x < 0 then |2x| = -2x. Substitute that into the equation |2x| + 3x = 10

OpenStudy (beth12345):

so is the answer x=-2 right for that part?

OpenStudy (beth12345):

and then make it become |2(-2)|+3(-2)=10?

OpenStudy (jamesj):

No. For x < 0 |2x| + 3x = 10 is equivalent to -2x + 3x = 10

OpenStudy (beth12345):

so how would i find what x is = to?

OpenStudy (jamesj):

C'mon. How do you solve that equation: -2x + 3x = 10

OpenStudy (jamesj):

-2x + 3x = x.

OpenStudy (beth12345):

oh, um oh i see, x =10?

OpenStudy (jamesj):

Yes. BUT, we started with x < 0 and have concluded that x = 10. That's a contradiction. Hence x = 10 is NOT a solution. Therefore the only solution is the one we found for x > 0.

OpenStudy (beth12345):

oh, ok thank you, i completely understand now :)